OpenStudy (rainbow_dash):

Jamila is making a triangular wall with building blocks. The top row has one block, the second row has three, the third has five, and so on. How many rows can she make with a set of 100 blocks?

OpenStudy (anonymous):

well this is arithmetic series

OpenStudy (anonymous):

what is a1, d , an ?

OpenStudy (rainbow_dash):

i thoguht it was pascal's triangle

OpenStudy (rainbow_dash):

or is pascals an arithmetic series?

OpenStudy (anonymous):

you can solve it using arithemetic series sum .. lets go with it .. a1 , d , an can you find them ?

OpenStudy (rainbow_dash):

an would be 100 yes? a1 would be 1 and d would be 2 i think.

OpenStudy (anonymous):

by an i mean general term

OpenStudy (anonymous):

an = a1 + (n-1)d

OpenStudy (rainbow_dash):

Oh, i'm clueless on an then :/

OpenStudy (rainbow_dash):

i'll just take numbers and plug them in though, thanks for equation :)

OpenStudy (anonymous):

d=2,a1=1 an = 1 + (n-1)*2 = 2n - 1 Sn = n * (an + a1) / 2 n * (2n - 1 + 1) / 2 = 100 200 = 2n^2 n = 10

OpenStudy (anonymous):

look at it

OpenStudy (anonymous):

Sn -> sum of n terms

OpenStudy (rainbow_dash):

so Sn= 1+(9)2..so 19?

OpenStudy (anonymous):

what i wrote up there is the whole solution.. first i found the general term formula (an) for example an(1) = a1 an(2) = a2 etc then i used Sn = n * (an + a1) / 2 this formula gives you the sum of "n" terms we only have 100 blocks so we need this sum to be 100 and we are looking for n -> number of rows plugging an,a1 into this formula gives you n = 10

OpenStudy (rainbow_dash):

oh dang lmao. my bad dude thanks :)

OpenStudy (anonymous):

yw;)